\begin{thebibliography}{10}

\bibitem{bolt}
E.~Bolthausen.
\newblock An estimate of the remainder in a combinatorial central limit
  theorem.
\newblock {\em Z. Wahrsch. Verw. Gebiete}, 66(3):379--386, 1984.

\bibitem{borel}
E.~Borel.
\newblock Sur les principes de la theorie cin\'etique des gaz.
\newblock {\em Annales de l'ecole normale sup.}, 23:9--32, 1906.

\bibitem{CM}
S.~Chatterjee and E.~Meckes.
\newblock {Multivariate central limit theorems via the method of exchangeable
  pairs}, 2005.
\newblock Preprint.

\bibitem{limit}
A.~D'Aristotile, P.~Diaconis, and C.~Newman.
\newblock Brownian motion and the classical groups.
\newblock In {\em Probability, Statistics and Their Applications: Papers in
  Honor of Rabi Bhattacharya}, volume~41 of {\em IMS Lecture Notes Monogr.
  Ser.}, pages 97--116. Inst. Math. Statist., Beachwood, OH, 2003.

\bibitem{diamal}
P.~Diaconis.
\newblock Application of the method of moments in probability and statistics.
\newblock In {\em Moments in Mathematics (San Antonio, Tex., 1987)}, volume~37
  of {\em Proc. Sympos. Appl. Math.}, pages 125--142. Amer. Math. Soc.,
  Providence, RI, 1987.

\bibitem{dia}
P.~Diaconis.
\newblock Patterns in eigenvalues: the 70th {J}osiah {W}illard {G}ibbs lecture.
\newblock {\em Bull. Amer. Math. Soc. (N.S.)}, 40(2):155--178 (electronic),
  2003.

\bibitem{del}
P.~Diaconis, M.~Eaton, and S.~Lauritzen.
\newblock Finite de {F}inetti theorems in linear models and multivariate
  analysis.
\newblock {\em Scand. J. Statist.}, 19(4):289--315, 1992.

\bibitem{diafree}
P.~Diaconis and D.~Freedman.
\newblock A dozen de {F}inetti-style results in search of a theory.
\newblock {\em Ann. Inst. H. Poincar\'e Probab. Statist.}, 23(2,
  suppl.):397--423, 1987.

\bibitem{diasha}
P.~Diaconis and M.~Shahshahani.
\newblock On the eigenvalues of random matrices.
\newblock {\em J. Appl. Probab.}, 31A:49--62, 1994.
\newblock Studies in applied probability.

\bibitem{hp}
F.~Hiai and D.~Petz.
\newblock {\em The Semicircle Law, Free Random Variables and Entropy},
  volume~77 of {\em Mathematical Surveys and Monographs}.
\newblock American Mathematical Society, Providence, RI, 2000.

\bibitem{hoeff}
W.~Hoeffding.
\newblock A combinatorial central limit theorem.
\newblock {\em Ann. Math. Statistics}, 22:558--566, 1951.

\bibitem{jia}
T.~Jiang.
\newblock Maxima of entries of {H}aar distributed matrices.
\newblock {\em Probab. Theory Related Fields}, 131(1):121--144, 2005.

\bibitem{joh}
K.~Johansson.
\newblock On random matrices from the compact classical groups.
\newblock {\em Ann. of Math. (2)}, 145(3):519--545, 1997.

\bibitem{steinbook}
C.~Stein.
\newblock {\em Approximate Computation of Expectations}.
\newblock Institute of Mathematical Statistics Lecture Notes---Monograph
  Series, 7. Institute of Mathematical Statistics, Hayward, CA, 1986.

\bibitem{steintech}
C.~Stein.
\newblock The accuracy of the normal approximation to the distribution of the
  traces of powers of random orthogonal matrices.
\newblock Technical Report No. 470, Stanford University Department of
  Statistics, 1995.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amsfonts,amssymb,amsthm,eucal,amsmath}
\usepackage{xy}
\input xy
\xyoption{all}
\allowdisplaybreaks
\newtheorem{thm}{Theorem}
\newtheorem{cor}[thm]{Corollary}
\newtheorem{lemma}[thm]{Lemma}

\newcommand{\R}{\mathbb{R}}
\newcommand{\C}{\mathbb{C}}

\newcommand{\E}{\mathbb{E}}
\newcommand{\N}{\mathbb{N}}

\newcommand{\ds}{\displaystyle}

\renewcommand{\P}{\mathbb{P}}

\renewcommand{\L}{\mathcal{L}}
\newcommand{\I}{\mathbb{I}}

\newcommand{\n}{\mathfrak{N}}
\renewcommand{\O}{\mathcal{O}}
\newcommand{\U}{\mathcal{U}}
\DeclareMathOperator{\tr}{Tr\,}

\renewcommand{\Re}{\operatorname{Re\,}}

\begin{document}

\title{Linear Functions on the Classical Matrix Groups}
\author[E.\ Meckes]{Elizabeth Meckes}\thanks{Research supported in
part by the ARCS Foundation. {\em AMS 2000 subject classifications.}
Primary 60F05; secondary 60B15, 60B10.  
{\em Key words and phrases.}
Stein's method, normal approximation, random matrices}

\maketitle

\begin{quote}
\begin{center}{\sc Abstract}\end{center}
{\small Let $M$ be a random matrix in the orthogonal group $\O_n$, distributed
according to Haar measure, and let $A$ be a fixed $n\times n$ matrix
over $\R$ such that $\tr(AA^t)=n$.  Then the total variation distance of
the random variable $\tr(AM)$ to a standard normal random variable 
is bounded by $\frac{2\sqrt{3}}
{n-1}$, and this rate is sharp up to the constant.  Analogous 
results are obtained for $M$ a random unitary 
matrix and $A$ a fixed $n\times n$ matrix over $\C$.  
The proofs are applications of a new abstract normal approximation theorem 
which extends Stein's method of
 exchangeable pairs to situations in which continuous symmetries are present.}

\end{quote}

\medskip


\section{Introduction}
Let $\O_n$ denote the group of $n\times n$ orthogonal matrices, and let 
$M$ be distributed according to Haar measure on $\O_n$.  
Let $A$ be a fixed $n\times n$ matrix over $\R$, subject to the condition
that $\tr(AA^t)=n$, and let $W=\tr(AM)$.  D'Aristotile, 
Diaconis, and Newman showed in \cite{limit} that 
$$\sup_{\substack{\tr(AA^t)=n
\\-\infty<x<\infty}}\left|\P(W\le x)-\Phi(x)\right|\to0$$
as $n\to\infty$.  Their argument uses classical methods involving 
sub-subsequences and tightness, and cannot be improved to yield
a theorem for finite $n$.  
Theorem \ref{orth-main1} below gives an explicit rate of 
convergence of the law of $W$ to the standard normal distribution
in the total variation
metric on probability measures, specifically,
\begin{equation}\label{mainst}
d(\L_W,\n(0,1))_{TV}\le\frac{2\sqrt{3}}{n-1}\end{equation}
for all $n\ge2.$

The history of this problem begins with the following theorem, first 
given rigorous proof by Borel in
\cite{borel}: let $X$ be a random vector on the unit sphere $S^{n-1}$, and
let $X_1$ be the first coordinate of $X$.  Then $\P\big(\sqrt{n}X_1\le t\big)
\longrightarrow\Phi(t)$ as $n\to\infty,$ where $\Phi(t)=\frac{1}{\sqrt{2\pi}}
\int_{-\infty}^te^{-\frac{x^2}{2}}dx.$  Since the first column of a 
Haar-distributed orthogonal matrix is uniformly distributed on the unit 
sphere, Borel's theorem follows from Theorem \ref{orth-main1} by taking 
$A=\sqrt{n}\oplus{\bf0}$.
Borel's theorem was generalized in one direction by Diaconis and 
Freedman \cite{diafree}, who proved the convergence of the first $k$ 
coordinates of $\sqrt{n}X$ to independent standard normal random variables
 in total variation
distance for $k=o(n)$; \cite{diafree} also contains a detailed history 
of this problem.  This line of research was further developed in
\cite{del}, where a total variation bound was given between an
$r\times r$ block of a random orthogonal matrix and an $r\times r$ 
matrix of independent standard Gaussians, for $r=O\left(n^{1/3}\right).$
This was later improved by Jiang (see \cite{jia}) to 
$r=O\left(n^{1/2}\right)$, which he proved was sharp.  
In the same paper, Jiang also showed that given a sequence of Haar distributed random matrices $\{M_n\}$, there is a sequence of Gaussian matrices $\{Y_n\}$ with $Y_j$ defined on the same probability space as $M_j$ such that if 
$$\epsilon_n=\max_{\substack{1\le i\le n\\ 1\le j\le m_n}}\left|\sqrt{n}M_{ij}-Y_{ij}\right|$$
with $m_n\le\frac{n}{\log^2n},$
then $\epsilon_n\to0$ in probability as $n\to\infty$.  Thus an 
$n\times\frac{n}{\log^2n}$ block of a Haar distributed matrix can be approximated by a Gaussian matrix `in probability'.
Theorem \ref{orth-main1}
 gives another
sense in which a random orthogonal matrix is close to a matrix of
independent normals by giving a uniform bound of distance to 
normal over all linear combinations of entries of $M$.

Another special case of Theorem \ref{orth-main1} is $A=I$, so that $W=\tr(M)$.  
Diaconis and Mallows (see \cite{diamal}) first proved that $\tr(M)$ is 
approximately normal; Stein \cite{steintech} and Johansson \cite{joh} later 
independently obtained fast 
rates of
convergence to normal of $\tr(M^k)$ for fixed $k$, with Johansson's rates 
an improvement on Stein's.  In studying eigenvalues of random orthogonal 
matrices, Diaconis and Shahshahani \cite{diasha} extended this to show that
the joint limiting distribution of $\tr(M), \tr(M^2),\ldots, \tr(M^k)$ 
converges to that of independent normal variables as $n\to\infty$, for
$k$ fixed.  

The other source of motivation for theorems like Theorem \ref{orth-main1} is
Hoeffding's combinatorial central limit theorem \cite{hoeff}, which can
be stated as follows.  Let $A=(a_{ij})$ be a fixed $n\times n$ 
matrix over $\R$, normalized to have row and column sums equal to zero and 
$\frac{1}{n-1}\sum_{i,j}a_{ij}^2=1$.  Let $\pi$ be a random permutation in
$S_n,$ and let $W(\pi)=\sum_i a_{i\pi(i)}.$  Then under certain conditions
on $A$, $W$ is approximately normal.  Later, Bolthausen \cite{bolt}
proved an explicit rate of convergence via Stein's method.  Note that if 
$$M_{ij}=\begin{cases}1&\pi(j)=i\\0&\mbox{ otherwise}\end{cases}$$
then $W=\tr(AM)$, and so Hoeffding's theorem is really a theorem about the
distribution of linear
functions on the set of permutation matrices.

The unitary group is another source of many important applications; see, e.g. 
\cite{dia}.  In Section \ref{unit}, the 
random variable $\tr(AM)$ for $A$ a fixed matrix over $\C$ and $M$ a random
unitary matrix distributed according to Haar measure on $\U_n$ is 
considered.  
The main theorem of the section, Theorem \ref{2main} gives a 
 bound on the total variation distance of 
$Re\big[\tr(AM)\big]$ to standard normal
 analogous to that of Theorem \ref{orth-main1}; this can be viewed as theorem 
about real-linear functions on $\U_n$.  Corollary \ref{complex} 
shows that in the limit, the 
complex random variable $\tr(AM)$ is close to standard complex normal.
The methods used here cannot be used directly to prove the convergence of
$\tr(AM)$ to the standard complex normal; they work for
approximation of real-valued random variables only.  
A version of the present methods in a multivariate
context is forthcoming in
\cite{CM}, which includes a rate of convergence for Corollary \ref{complex}.

{\bf Notation and Conventions.}
The total variation distance $d_{TV}(\mu,\nu)$ between the measures $\mu$ and
$\nu$ on $\R$ is defined by
$$d_{TV}(\mu,\nu)=\sup_A\big|\mu(A)-\nu(A)|,$$
where the supremum is over measurable sets $A$.  This is equivalent to
$$d_{TV}(\mu,\nu)=\frac{1}{2}\sup_{f}\left|\int f(t)d\mu(t)-
\int f(t)d\nu(t)\right|,$$
where the supremum is taken over continuous functions which are bounded
by 1 and vanish at infinity; this is the definition used in what follows. 
The total variation distance between two random variables $X$ and $Y$ is
defined to be the total variation distance between their distributions:
$$d_{TV}(X,Y)=\sup_A\big|\P(X\in A)-\P(Y\in A)\big|=\frac{1}{2}\sup_{f}
\big|\E f(X)-\E f(Y)\big|.$$
We will use $\n(\mu,\sigma^2)$ to denote the normal distribution on $\R$ with 
mean $\mu$ and variance $\sigma^2$.  

\medskip

{\bf Acknowledgements.}
I would like to thank Persi Diaconis for sharing his many
insights with me.

\section{An abstract normal approximation theorem}\label{chap-abs1}
In this section, a general approach for normal approximation to random
variables with continuous symmetries is developed.  The ideas which give 
rise to Theorem \ref{abscont} below first appeared in Stein \cite{steintech},
where fast rates of convergence to Gaussian (as $n\to\infty$) 
were obtained for $\tr(M^k)$,
with $k\in\N$ fixed and $M$ a random $n\times n$ orthogonal
matrix.  
\begin{thm}\label{abscont}
Suppose that $(W,W_\epsilon)$ is a family of exchangeable pairs defined
on a common probability space with $\E W=0$ and $\E W^2=\sigma^2$.
Suppose that there are functions $\alpha$ and $\beta$ with
$$\E|\alpha(\sigma^{-1}W)|<\infty,\qquad\E|\beta(\sigma^{-1}W)|<\infty,$$
and a constant $\lambda$ such that 
\begin{enumerate}
\item $$\frac{1}{\epsilon^2}\E\left[W_\epsilon-W\big|
W\right]=-\lambda W+o(1)\alpha(W),$$ \label{lin-cond}
\item $$\frac{1}{\epsilon^2}\E\left[(W_\epsilon-W)^2\big|
W\right]=2\lambda\sigma^2 +E\sigma^2+o(1)\beta(W),$$ \label{quad-cond}
\item $$\frac{1}{\epsilon^2}\E\left|W_\epsilon-W\right|^3
=o(1),$$ \label{cube-cond}
\end{enumerate}
where $o(1)$ refers to the limit as $\epsilon\to0$, with 
the implied constants deterministic.

Then
$$d_{T.V.}(W,Z)\le \frac{1}{\lambda}\E\left|E\right|,$$
where $Z\sim\n(0,\sigma^2)$.
\end{thm}

\medskip

{\bf Remark:} The factor of $\frac{1}{\epsilon^2}$ in each of 
the three expressions above could be replaced by a general function 
$f(\epsilon)$.  In practice, $W_\epsilon$ is typically constructed such that
$W_\epsilon-W=O(\epsilon)$.  This makes it clear that
$f(\epsilon)=\frac{1}{\epsilon^2}$ is
the suitable choice for condition \ref{quad-cond}.  It is less clear that
$f(\epsilon)=\frac{1}{\epsilon^2}$ is the suitable choice for 
condition \ref{lin-cond}.  In the applications given here, while 
$W_\epsilon-W=O(\epsilon)$, symmetry conditions imply that
$$\E\left[W_\epsilon-W\big|W\right]=O(\epsilon^2).$$

\bigskip

Before beginning the proof, some background on Stein's method is helpful.
The following lemma is key.
\begin{lemma}[Stein]\label{gauss}Let $Z\sim\n(0,1).$ 
Then
\begin{enumerate}
\item For all $f\in C^1_o(\R),$
$$\E\big[f'(Z)-Zf(Z)\big]=0.$$
\item \label{converse}If $Y$ is a random variable such that 
$$\E\big[f'(Y)-Yf(Y)\big]=0$$
for all $f\in C_b^1(\R),$ then $\L(Y)=\L(Z);$ i.e., $Y$ is also distributed
as a standard Gaussian random variable.
\item For $g:\R\to\R$ with $\E g(Z)<\infty$ given, the function
\begin{equation}\label{U_odef}
U_og(t)=e^{t^2/2}\int_{-\infty}^t\big[g(x)-\E g(Z)\big]e^{-x^2/2}dx.
\end{equation}
is a solution to the differential equation
$$f'(x)-xf(x)=g(x)-\E g(Z).$$
\end{enumerate}
\end{lemma}

\medskip

The lemma says that the standard Gaussian distribution $\gamma$ on $\R$ is the 
unique distribution with the property that $\int_{\R}(f'(x)-xf(x))d\gamma(x)$ 
is always zero.  The idea of Stein's method is that
if $W$ is a random variable such that $\E\big[f'(W)-Wf(W)\big]$ is always
{\em small}, then the distribution of $W$ is close $\gamma$.  
There are several 
approaches to bounding this quantity; the approach taken here is modelled 
on the method of exchangeable pairs (see \cite{steinbook}).
In any of the approaches, the following bounds on $U_o$ are useful.

\begin{lemma}[Stein]\label{stbds}
Let $U_o$ be the operator defined in equation (\ref{U_odef}).  Then
\begin{enumerate}
\item $\|U_og\|_\infty\le \sqrt{\frac{\pi}{2}}\|g-\E g(Z)\|_\infty
\le \sqrt{2\pi}\|g\|_\infty$
\item $\|(U_og)'\|_\infty\le 2\|g-\E g(Z)\|_\infty\le 4\|g\|_\infty$
\item $\|(U_og)''\|_\infty\le 2\|g'\|_\infty$
\end{enumerate}
\end{lemma}

\medskip

With this background, the proof of Theorem \ref{abscont} is straightforward.
\begin{proof}[Proof of Theorem \ref{abscont}]
By considering $\sigma^{-1}W$ instead of $W$, we may without loss assume 
that $\sigma=1$.  
For $g\in C^\infty_o(\R)$ fixed, let $f$ be the solution given in equation 
\eqref{U_odef} to the differential 
equation
$$f'(x)-xf(x)=g(x)-\E g(Z).$$
Fix $\epsilon$.  By the exchangeability of $(W,W_\epsilon)$, 
\begin{equation}\label{exch}\begin{split}
0&=\E\left[(W_\epsilon-W)(f(W_\epsilon)+f(W))\right]\\
&=\E\left[(W_\epsilon-W)(f(W_\epsilon)-f(W))+2(W_\epsilon-W)f(W)\right]\\
&=\E\left[\E\left[(W_\epsilon-W)^2\big|W\right]f'(W)+2\E\left[(W_\epsilon-
W)\big|W\right]f(W)+R\right],
\end{split}\end{equation}
where $R$ is the error in the derivative approximation.  By Taylor's theorem
and Lemma \ref{stbds},
$$|R|\le\frac{\|f''\|_\infty}{2}|W_\epsilon-W|^3\le\|g'\|_\infty|W_\epsilon-W
|^3,$$ and so 
$$\lim_{\epsilon\to0}\frac{1}{\epsilon^2}\E|R|=0.$$
Dividing both sides of (\ref{exch}) by $2\lambda\epsilon^2$ and 
taking the limit 
as $\epsilon\to0$ gives:
$$0=\E\left[f'(W)-Wf(W)+\frac{E}{2\lambda}f'(W)\right]=\E\left[g(W)-g(Z)+
\frac{E}{2\lambda}f'(W)\right].$$
Rearranging and applying the bound on $\|f'\|$ 
from Lemma
\ref{stbds} yields
\begin{equation*}
\big|\E g(W)-\E g(Z)\big|\le\frac{2\|g\|_\infty}{\lambda}\E|E|.
\end{equation*}
Since $C^\infty_o(\R)$ is dense (with respect to the supremum norm)
in the class of bounded 
continuous functions vanishing at infinity, this completes the proof.
\end{proof}

\section{The Orthogonal Group}\label{orth}
This section is mainly devoted to the proof of the following theorem.

\begin{thm}\label{orth-main1}
Let $A$ be a fixed $n\times n$ matrix over $\R$ such that $\tr(AA^t)=n,$
 $M\in\O_n$ distributed according to Haar measure, and $W=\tr(AM)$.  Let $Z$
be a standard normal random variable. 
Then for $n>1$,
\begin{equation}\label{thm}
d(W,Z)_{TV}\le\frac{2\sqrt{3}}{n-1}.\end{equation}
\end{thm}
\vspace{5mm}
The bound in Theorem \ref{orth-main1} is sharp up to the constant; consider the
matrix $A=\sqrt{n}\oplus{\bf0}$ where ${\bf0}$ is the $n-1\times n-1$
matrix with all zeros.  For this $A$, Theorem \ref{orth-main1} reproves the following
theorem, proved in \cite{diafree} with slightly worse constant
\begin{thm}\label{sphere}
Let ${\bf x} \in\sqrt{n}S^{n-1}$ be uniformly distributed, and let $Z$ be
a standard normal random variable.  Then
$$d_{TV}(x_1,Z)\le\frac{2\sqrt{3}}{n-1}.$$
\end{thm}

It is shown in \cite{diafree} that the order of this error term is correct.

\begin{proof}[Proof of Theorem \ref{orth-main1}]
First note that one can
assume without loss of generality that $A$ is diagonal: let $A=UDV$ be the
singular value decomposition of $A$.  Then $W=\tr(UDVM)=\tr(DVMU)$, and the
distribution of $VMU$ is the same as the distribution of $M$ by the 
translation invariance of Haar measure.  

Now define the pair $(W,W_\epsilon)$ for
each $\epsilon$ as follows.  Choose $H=(h_{ij})\in\O(n)$ according to Haar 
measure, 
independent of $M$, and let $M_\epsilon=HA_\epsilon H^tM$, where
$$A_\epsilon=\begin{bmatrix}\sqrt{1-\epsilon^2}&\epsilon&&&\\-\epsilon&\sqrt
{1-\epsilon^2}&&0&\\&&1&&\\&0&&\ddots&\\&&&&1\end{bmatrix},$$
thus $M_\epsilon$ can be thought of as a small random rotation of $M$.  
Let $W_\epsilon=W(M_\epsilon)$; $(W,W_\epsilon)$ is an exchangeable pair
by construction.

It is convenient to rewrite $M_\epsilon$ as follows.  Let 
$I_2$ be the $2\times2$ identity matrix, $K$ the $n\times
2$ matrix consisting of the first two columns of $H$, and let 
$$C_2=\begin{bmatrix}0&1\\-1&0\end{bmatrix}.$$  Then 
\begin{equation*}\begin{split}
M_\epsilon&=M+K\big[(\sqrt{1-\epsilon^2}-1)I_2+\epsilon C_2\big]K^tM
\\&=M+K\left[\left(-\frac{\epsilon^2}{2}+O(\epsilon^4)\right)I_2+
\epsilon C_2\right]K^tM,\end{split}\end{equation*}
and so 
\begin{equation}
W_\epsilon-W=\epsilon\left[\left(-\frac{
\epsilon}{2}+O(\epsilon^3)\right)\tr(AKK^tM)+\tr(AKC_2K^tM)\right]
\label{diff}\end{equation}

Now, the distribution of $H$ is unchanged by multiplying a fixed 
row or column by $-1$ and $H$ is orthogonal, 
thus $\E h_{ij}h_{k\ell}=\frac{1}{n}\delta_{ik}
\delta_{j\ell}$.  This implies that 
$$\E \big[KK^t\big]=\frac{2}{n}I_n$$
and $$\E \big[KCK^t\big]=0;$$
combining this with
 (\ref{diff}) yields:
\begin{equation}\begin{split}
\frac{n}{\epsilon^2}\E&\left[(W_\epsilon-W)\big|W\right]\\&=
-\frac{n}{2}\E\left[\E\left[\tr(AKK^tM)\big|M\right]\big|W\right]+
\frac{n}{\epsilon}\E\left[\E\left[\tr(AKC_2K^tM)\big|M
\right]\big|W\right]+O(\epsilon)\\&=-\E\left[\E\left[\tr(AM)\big|M\right]
\big|W\right]+O(\epsilon)
\nonumber\\&=-W+O(\epsilon),
\end{split}\end{equation}
where the independence of $M$ and $H$ has been used to get the third line,
and the implied constants in the $O(\epsilon)$ here and in what follows  
may depend on $n$.
Condition \ref{lin-cond} of Theorem \ref{abscont} is thus 
satisfied with $\lambda=\frac{1}{n}$.

Recall now that $A$ is assumed to be diagonal.  
The second condition of Theorem \ref{abscont} can also be verified using
 the expression in (\ref{diff}) as follows.
\begin{equation}\begin{split}
\frac{n}{2\epsilon^2}&\E\left[(W_\epsilon-W)^2\big|W\right]\\
&=\frac{n}{2}\E\left[\E\left[(\tr(AKC_2K^tM))^2\big|M\right]\big|
W\right]+O(\epsilon)\\
&=\left.\frac{n}{2}\E\left[\sum_{i,j}\sum_{\substack{i'\neq i\\
j'\neq j}}m_{i'i}m_{j'j}a_{ii}a_{jj}\label{secdiff1}
\E\left[(h_{i1}h_{i'2}-h_{i2}h_{i'
1})(h_{j1}h_{j'2}-h_{j2}h_{j'1})\big|M\right]\right|W\right]+O(\epsilon),
\end{split}\end{equation} 
where the conditions on $i'$ and $j'$ are justified as the expression inside
the expectation is identically zero when either $i=i'$ or $j=j'$.  

Standard techniques are available for computing the mixed moments of 
entries of $H$; see e.g. \cite{hp}, section 4.2.  Using these techniques
 and the independence of $M$ and $H$ gives that
for $i'\neq i$ and $j'\neq j$,
\begin{equation}\label{useful-fourth}
\E\left[(h_{i1}h_{i'2}-h_{i2}h_{i'1})(h_{j1}h_{j'2}-h_{j2}h_{j'1})\big|M
\right]=\frac{2}{n(n-1)}\Big[\delta_{ij}\delta_{i'j'}-\delta_{ij'}
\delta_{ji'}\Big];\end{equation}
putting this into (\ref{secdiff1}) yields
\begin{eqnarray*}\nonumber
\frac{n}{2\epsilon^2}\E\left[(W_\epsilon-W)^2\big|W\right]
&=&\frac{1}{n-1}\sum_{i,j}\sum_{\substack{i'\neq i\\j'\neq j}}m_{i'i}m_{j'j}
a_{ii}a_{jj}\Big[\delta_{i'j'}\delta_{ij}-\delta_{ij'}\delta_{ji'}\Big]
+O(\epsilon)\\
\nonumber&=&\frac{1}{n-1}\left[\sum_ia_{ii}^2\left[(M^tM)_{ii}-m_{ii}^2\right]
-\sum_{i,i'\neq i}(MA)_{ii'}(MA)_{i'i}\right]+O(\epsilon)\\\nonumber&=&
\frac{1}{n-1}\left[n-\sum_ia_{ii}^2m_{ii}^2-\left[\tr((MA)^2)-\sum_ia_{ii}^2
m_{ii}^2\right]\right]+O(\epsilon)
\\&=&1+\frac{1}{n-1}
\left[1-\tr((AM)^2)\right]+O(\epsilon),
\end{eqnarray*}
thus 
\begin{equation}\label{secdiff2}
\lim_{\epsilon\to0}\frac{1}{\epsilon^2}\E\left[(W_\epsilon-W)^2\big|W\right]
=\frac{2}{n}+\frac{2}{n(n-1)}\left[1-\tr((AM)^2)\right]\end{equation}
and so 
\begin{equation}\label{orthE}E=\frac{2}{n(n-1)}\left[1-\tr((AM)^2)\right].
\end{equation}

Finally, \eqref{diff} gives immediately that 
$$\E\left[|W_\epsilon-W|^3\big|W\right]=O(\epsilon^3).$$
It remains to bound $n\E|E|.$

\begin{eqnarray}
\E\left[\tr((AM)^2)\right]&=&\E\left[\sum_{i,j}a_{ii}a_{jj}m_{ij}m_{ji}\right]
\nonumber\\
&=&\frac{1}{n}\sum_ia_{ii}^2=1\label{mean},
\end{eqnarray}
and
\begin{equation*}\begin{split}
\E\big[(\tr(&(AM)^2))^2\big]\\&=\E\left[\left(\sum_{i,j}a_{ii}a_{jj}m_{ij}
m_{ji}\right)\left(\sum_{k,l}a_{kk}a_{ll}m_{kl}m_{lk}\right)\right]\nonumber
\\\nonumber&=
\sum_{i,j,k,l}a_{ii}a_{jj}a_{kk}a_{ll}\left[\frac{n+1}{(n-1)n(n+2)}\left[
\delta_{ij}\delta_{kl}\big(1-\delta_{ik}\big)+
\delta_{ik}\delta_{jl}\big(1-\delta_{ij}\big)+\right.\right.
\\\nonumber&\left.\left.\qquad\qquad\qquad\qquad\qquad
\delta_{il}\delta_{jk}\big(1-\delta_{ij}\big)
\right]+\frac{3}{n(n+2)}\I(i=j=k=l)\right]\\
&=\frac{n+1}{n(n-1)(n+2)}\left(\sideset{}{'}\sum_{i,k}a_{ii}^2
a_{kk}^2+\sideset{}{'}\sum_{i,j}a_{ii}^2a_{jj}^2+\sideset{}{'}\sum_{
i,j}a_{ii}^2a_{jj}^2\right)+\frac{3}{n(n+2)}\sum_{i}a_{ii}^4\nonumber
\end{split}\end{equation*}
Now,
\begin{equation*}\begin{split}
\sideset{}{'}\sum_{i,j}a_{ii}^2a_{jj}^2&=\sum_ia_{ii}^2(n-a_{ii}^2)
=n^2-\sum_ia_{ii}^4.
\end{split}\end{equation*}
Applying this above gives
\begin{equation}\begin{split}
\E\big[(\tr((AM)^2))^2\big]&
=\frac{3(n+1)n^2}{(n-1)n(n+2)}-
\frac{3(n+1)}{(n-1)n(n+2)}\sum_ia_{ii}^4+\frac{3}{n(n+2)}\sum_ia_{ii}^4
\\&\le3+\frac{6}{(n-1)(n+2)}.\label{4th}
\end{split}\end{equation}

Putting these estimates into Theorem \ref{abscont} gives:
\begin{equation}d_{T.V.}(W,Z)\le\frac{2\sqrt{2+\frac{6}{(n-1)(n+2)}}}{(n-1)}.
\end{equation}
Noting that $\frac{6}{(n-1)(n+2)}\le1$ for
$n\ge3$ and that the bound in Theorem \ref{orth-main1} is trivially true for 
$n=2$ completes the proof.  
\end{proof}

\section{The Unitary Group}\label{unit}
Now let $M\in\U_n$ be distributed according 
to Haar measure, $A$ be an $n\times n$ matrix over $\C$, and
$W=\tr(AM)$.  In \cite{limit} it was shown that if $M=\Gamma+i\Lambda$ and 
$A$ and $B$ are fixed real diagonal matrices with $\tr (AA^t)=\tr(BB^t)=n,$
then $\tr(A\Gamma)+i\tr(B\Lambda)$ converges in distribution to a standard
complex normal random variable.  
This implies in particular that $\Re(W)$ converges in
distribution to $\n\left(0,\frac{1}{2}\right).$  The main theorem of 
this section gives a rate of this convergence in total 
variation distance.  

A more natural question might be the convergence of 
$W$ to a standard complex random variable.  As this is a multivariate problem,
Theorem \ref{abscont} cannot be applied.  A multivariate version of 
Theorem \ref{abscont}
is forthcoming in \cite{CM}, which also includes a rate of 
convergence of $W$ to 
a standard complex Gaussian random variable.

\begin{thm}\label{2main}
With $M$, $A$, and $W$ as above, let $W_\theta$ be the inner product of
$W$ with the unit vector making angle $\theta$ with the real axis.  Then
\begin{equation}
d_{TV}\left(W_\theta,\n\left(0,\frac{1}{2}\right)\right)\le\frac{c}{n}
\end{equation}
for a constant $c$ which is independent of $\theta$.
\end{thm}

The constant $c$ is asymptotically equal to $2\sqrt{2}$; for $n\ge 8$ it can
be taken to be 4.

\medskip

\begin{proof}
To prove the theorem, first note that it suffices to consider the case 
$\theta=0$, that is, to prove that 
$$d_{TV}\left(\Re(W),\n\left(0,\frac{1}{2}\right)\right)\le\frac{c}{n}.$$
The theorem then follows as stated since the distribution of $W$ is 
invariant under multiplication by any complex number of unit modulus.  Also,
$A$ can again be assumed diagonal with positive real entries by the singular 
value decomposition.

The proof is almost identical to the orthogonal case.  
Let $H\in\U_n$ be a random unitary matrix, independent of 
$M$, and let $M_\epsilon=HA_\epsilon H^*M$, where
$$A_\epsilon=\begin{bmatrix}\sqrt{1-\epsilon^2}&\epsilon&&&\\-\epsilon&\sqrt
{1-\epsilon^2}&&0&\\&&1&&\\&0&&\ddots&\\&&&&1\end{bmatrix}.$$
Let $W_\epsilon=W(M_
\epsilon)$.

Let
$I_2$ be the $2\times2$ identity matrix, $K$ the $n\times
2$ matrix consisting of the first two columns of $H$, and let 
$$C_2=\begin{bmatrix}0&1\\-1&0\end{bmatrix}.$$  Then
\begin{eqnarray}
W_\epsilon-W&=&\tr\left(\left(-\frac{\epsilon^2}{2}+O(\epsilon^4)\right)AKK^*M+
\epsilon AKC_2K^*M\right)\nonumber\\&=&\epsilon\left[\left(-\frac{
\epsilon}{2}+O(\epsilon^3)\right)\tr(AKK^*M)+\tr(AKC_2K^*M)\right].
\label{diff'}\end{eqnarray}

Let $W^r=Re(W)$ and $W^r_\epsilon=Re(W_\epsilon)$.  
As in the orthogonal case, to verify the conditions of Theorem \ref{abscont}
various mixed moments of the entries of $H$ are needed.  The relevant
unitary integrals can also be found in \cite{hp}, section 4.2.  They imply 
in particular that
\begin{eqnarray}
\E\left[KK^*_{ij}\right]&=&\frac{2}{n}\delta_{ij}\label{KK^*}\\
\E\left[KC_2K^*_{ij}\right]&=&0\label{KCK},
\end{eqnarray}
thus
\begin{equation}\label{firstdiff'}
\lim_{\epsilon\to0}\frac{n}{\epsilon^2}\E\left[W^r_\epsilon-W^r\big|W\right]=
-W^r;\end{equation}
condition \ref{lin-cond} is satisfied with $\lambda=\frac{1}{n}.$
Also by (\ref{diff'}),
\begin{equation}\begin{split}
\lim_{\epsilon\to0}\frac{n}{2\epsilon^2}\E[(W_\epsilon^r &
-W^r)^2\big|W]\\&=\lim_{\epsilon\to0}\frac{n}{2}\E\left[\left
(Re(\tr(AKC_2K^*M))\right)^2\big|W\right]\\
&=\frac{n}{4}Re\,\E\left[\sum_{i,j,k,l}a_{ii}m_{ji}a_{kk}m_{lk}(h_{i1}
\overline{h}_{j2}-h_{i2}\overline{h}_{j1})(h_{k1}
\overline{h}_{l2}-h_{k2}\overline{h}_{l1})\right.+
\\\label{secdiff1'}&\left.\left.\quad\qquad\qquad
\phantom{\sum_{i,j,k,l}}a_{ii}m_{ji}a_{kk}\overline{m}_{lk}(h_{i1}
\overline{h}_{j2}-h_{i2}\overline{h}_{j1})(\overline{h}_{k1}
h_{l2}-\overline{h}_{k2}h_{l1})\right|W\right]
\end{split}\end{equation}
Using the formulae from \cite{hp}, it is straightforward to show that
\begin{equation}\begin{split}\label{prod1'}
\E[(h_{i1}\overline{h}_{j2}-h_{i2}\overline{h}
_{j1})&(h_{k1}\overline{h}_{l2}-h_{k2}\overline{h}_{l1})]\\&=-\frac{2\delta_{il}
\delta_{jk}(1-\delta_{ij})}{(n-1)(n+1)}+
\frac{2\delta_{ij}\delta_{k\ell}(1-\delta_{ik})}{(n-1)n(n+1)}
-\frac{2\I(i=j=k=l)}{n(n+1)}
\end{split}\end{equation}
and
\begin{eqnarray}
\E\big[(h_{i1}\overline{h}_{j2}&&\hspace{-9mm}-\hspace{1mm}h_{i2}
\overline{h}_{j1})(\overline{h}_{k1}h_{l2}-\overline{h}_{k2}h_{l1})\big]
\hspace{3in}
\nonumber\\
&=&\frac{2\left(\delta_{ik}\delta_{jl}(1-\delta_{ij})
\right)}{(n-1)(n+1)}-\frac{2\left(\delta_{ij}\delta_{kl}(1-\delta_{ik})\right)}
{n(n-1)(n+1)}\label{prod2}+\frac{2\I(i=j=k=l)}{n(n+1)}.
\end{eqnarray}

Let $\ds\sideset{}{'}\sum_{i,j}$ stand for summing over all pairs $(i,j)$ where
$i$ and $j$ are distinct.  Putting (\ref{prod1'}) and (\ref{prod2}) into 
(\ref{secdiff1'}) and using the
independence of $M$ and $H$ gives:

\begin{equation*}\begin{split}
\lim_{\epsilon\to0}\frac{n}{2\epsilon^2}
\E\big[\big(&W_\epsilon^r-W^r\big)
^2\big|W\big]\\
&=\frac{n}{2(n-1)(n+1)}Re\,\E\left[\sum_{i,j,k,\ell}a_{ii}m_{ji}
a_{kk}m_{\ell k}\left(-\delta_{i\ell}\delta_{jk}(1-\delta_{ij})+
\frac{1}{n}\delta_{ij}\delta_{k\ell}(1-\delta_{ik})\right.\right.\\
&\hspace{3.5in}-\left.\left(\frac{n-1}{n}\right)\I(i=j=k=\ell)\right)
\\&\hspace{1.75in}
+\sum_{i,j,k,\ell}a_{ii}m_{ji}a_{kk}m_{\ell k}
\left(\delta_{ik}\delta_{j\ell}(1-\delta_{ij})-\frac{1}{n}\delta_{ij}
\delta_{k\ell}(1-\delta_{ik})\right.\\&\hspace{3.25in}
\left.\left.\phantom{\sum_i}\left.+\left(\frac{n-1}{n}\right)\I(i=j=k=l)
\right)
\right|W\right]
\\&=\frac{n}{2(n-1)(n+1)}Re\,\E\left[-\sideset{}{'}
\sum_{i,j}a_{ii}a_{jj}m_{ij}m_{ji}\right.+\frac{1}{n}
\sideset{}{'}\sum_{i,k}a_{ii}a_{kk}m_{ii}m_{kk}
\\&\hspace{2in}-\left(\frac{n-1}{n}\right)\sum_ia_{ii}^2m_{ii}^2+
\sideset{}{'}\sum_{i,j}a_{ii}^2|m_{ji}|^2
\\&\hspace{2in}-\frac{1}{n}\sideset{}{'}\sum_{i,k}a_{ii}a
_{kk}m_{ii}\overline{m}_{kk}\left.\left.+
\frac{n-1}{n}\sum_ia_{ii}^2|m_{ii}|^2\right|W\right]\\
&=\frac{n}{2(n-1)(n+1)}Re\,\E\left[-\left(\tr((AM)^2)-\sum_i(AM)_{ii}
^2\right)\right.+\frac{1}{n}\left(W^2-
\sum_i(AM)_{ii}^2\right)\\&\hspace{2in}
-\left(\frac{n-1}{n}\right)
\sum_i(AM)_{ii}^2+
\sum_ia_{ii}^2(1-|m_{ii}|^2)\\&\hspace{2in}\left.\left.
-\frac{1}{n}\left(\left|W\right|^2-\sum_ia_{ii}^2|m_{ii}|^2\right)
+\frac{n-1}{n}\sum_i
a_{ii}^2|m_{ii}|^2\right|W\right]
\\&=
\frac{1}{2}+\frac{1}{2(n-1)(n+1)}\\&
\left.+\frac{n}{2(n-1)(n+1)}Re\,\E\left[-\tr((AM)^2)
+\frac{W^2-|W|^2}{n}\right|W\right].
\end{split}\end{equation*}
Condition (2) of Theorem \ref{abscont} is thus satisfied with 
\begin{equation}\begin{split}\label{unitE}
nE&=\frac{1}{2(n-1)(n+1)}+\left.\frac{n}{2(n-1)(n+1)}
Re\,\E\left[-\tr((AM)^2)
+\frac{W^2-|W|^2}{n}\right|W\right].
\end{split}\end{equation}
It remains to estimate $n\E|E|$.  First,
\begin{eqnarray*}
\E\left|\tr((AM)^2)\right|&=&\E\sqrt{\sum_{i,j,k,l}a_{ii}a_{jj}m_{ij}m_{ji}
a_{kk}a_{ll}\overline{m}_{kl}\overline{m}_{lk}}\\
&\le&\sqrt{\sum_{i,j,k,l}a_{ii}a_{jj}a_{kk}a_{ll}
\E\left[m_{ij}m_{ji}\overline{m}_{kl}\overline{m}_{lk}\right]}\\
&=&\sqrt{\frac{2n^2}{(n-1)(n+1)}-\frac{2}{(n-1)n(n+1)}\left(\sum_i
a_{ii}^4\right)}\\&\le&\sqrt{2+\frac{1}{n^2-1}},
\end{eqnarray*}
using the formulae of \cite{hp} to evaluate the integrals.

Next,
\begin{eqnarray*}
\E|W|^2&=&\E\left[\sum_{i,j}a_{ii}a_{jj}m_{ii}\overline{m}_{jj}
\right]\\&=&\frac{1}{n}\sum_ia_{ii}^2\\&=&1.
\end{eqnarray*}

Putting these estimates into (\ref{unitE}) proves the theorem.

\end{proof}

\medskip

Theorem \ref{2main} yields the following bivariate corollary, which 
can also be seen as a corollary of the main unitary lemma of \cite{limit}.

\begin{cor}\label{complex}
Let $M$ be a random unitary matrix, $A$ a fixed $n\times n$ matrix over
$\C$ with $\tr(AA^*)=n$, and let $W=\tr(AM)$.  Then the distribution
of $W$ converges to the standard complex normal distribution in the weak-star
topology.
\end{cor}

\begin{proof}
The result follows immediately from Theorem \ref{2main} by considering 
the characteristic function of $W$.
\end{proof}

\bibliographystyle{plain}
\bibliography{matrices2}

\bigskip

\noindent Elizabeth S. Meckes\\
Department of Mathematics\\
Stanford University\\
Stanford, CA 94305\\

\noindent meckes@math.stanford.edu\\

\end{document}